plz help me solve this: Gp has 6 and 24 as its 2nd and 4rh terms respectively, now I need the first four terms
Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

an=a1*r^(n-1)
6=a1*r; r=6/a1
24=a1*r^3
substitute
24=a1*216/a1^3=216/a1^2
a1^2=216/24=9
a1=3
r=2
The first four terms are 3,6,12, and 24.

Answer by MathTherapy(10552)


plz help me solve this: Gp has 6 and 24 as its 2nd and 4rh terms respectively, now I need the first four terms
You already have the 2nd and 4th. The 1st and 3rd can be: